OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through (2, –1) and has a slope of –5. A. (y – 1) = –5(x + 2) B. (y – 2) = –5(x – 1) C. (y + 1) = –5(x – 2) D. (y + 2) = –5(x – 1)

OpenStudy (butterflydreamer):

this one is just like the other one :) Use the point -slope formula: y - y1 = m ( x - x1) We are given slope (m) = -5 and that it passes through ( 2 , -1) ---> this means x1=2 , y1 = -1 So substitute in m = -5 , x1 = 2 and y1 = -1 into the point-form formula :)

OpenStudy (anonymous):

so would it be D?

OpenStudy (butterflydreamer):